1. A lampshade is in the form of a frustum of a cone of which the top and the bottom radius is thrice the top radius, and the height is 14cm.

The volume of the lampshade is 763cm³


Find:

a. The area of material required to cover the lampshade

b. Bottom Radius ?


1
Expert's answer
2020-07-08T18:35:08-0400

Solution 1. a

Let the volume of the frustum be V, the top radius be r and bottom radius be R and vertical height be h.

To calculate the volume V, we use the formula

"V=\u03c0h\/3(R^2+Rr+r^2)"


We are given that V=763cm3 , h=14cm, R=3r

We then substitute the figures in the formula as follows;

763=14π/3[ (3r)2 + (3r)(r)+r2]

763=14π/3[9r2+3r2+r2], we sum the terms in the brackets.

763=14π/3[13r2] then we multiply 14 by 13

763=182πr2/3 When we multiply by 3/182π both sides, we get the following:

763(3/182π)= 182πr2/3 (3/182π), we are left with r2 on the right hand side giving

2289/182π=r2 To find r alone, we need to take the square root of both sides

"\sqrt {(2289/182π)}= \sqrt{r^2}"

Working out the square root of both sides we get r=2.00cm

Therefore, radius r=2.00cm and R= 3r=3(2)=6.00cm

We want to find the area of material required to cover the lampshade. This is basically the surface area and it is given by;


"Surface Area=\u03c0l(R+r)" But "l" is the slant height and is given by

"l= \sqrt{h^2+R^2}"

h=14, R=6

"l=\sqrt{(14^2+6^2 )}"

"l=\sqrt{(196+36)}"

"l=\sqrt{232}"

l=15.23cm

Hence Surface Area=15.23π(6+3)

=15.23π(9)

430.62cm2


Solution 1. b

Bottom radius R =3r

= 3(2.00)

=6.00cm